OpenStudy (anonymous):

Suppose f ( x, y ) = [x/y], P = ( −4, −1 ) and v = −2 I − 2 J. A. Find the gradient of f. ∇f = ??? I + ??? J Note: Your answers should be expressions of x and y; e.g. "3x - 4y" B. Find the gradient of f at the point P. ( ∇f ) ( P ) = ??? I + ??? J C. Find the directional derivative of f at P in the direction of v, where the direction u of a vector v is the unit vector obtained by normalizing that vector, i.e., u = [(v )/(|| v || )]. Du f = ? D. Find the maximum rate of change of f at P. ? E. Find the (unit) direction vector in which the maximum rate of change occurs at P. u = ? i + ? j

OpenStudy (turingtest):

A: f(x,y)=x/y partial w/respect to x is fx(x,y)=1/y i partial w/respect to y is fy(x,y)=-1/(2y^2) j (Del)f=(1/y)i-1/(2y^2)j try B yourself first from here...